OpenStudy (anonymous):

1/3+1/6+1/12+1/24..........................infinite=?

OpenStudy (anonymous):

\[[\sqrt{(3n)^3} -2]/(9n-3)\]

OpenStudy (anonymous):

its wrong

OpenStudy (anonymous):

\[\frac{ 1 }{ 3}+\frac{ 1 }{ 6 }+\frac{ 1 }{ 12 }+\frac{ 1 }{ 24 }+.....\] \[\frac{ 1 }{ 3}+\frac{ 1 }{ 3\times2 }+\frac{ 1 }{ 3\times4 }+\frac{ 1 }{ 3\times8 }+.....\] \[\frac{ 1 }{ 3}(1+\frac{ 1 }{ 2 }+\frac{ 1 }{ 4 }+\frac{ 1 }{ 8 }+.....)\] \[\frac{ 1 }{ 3}(\frac{ 1 }{2^{0} }+\frac{ 1 }{ 2^{1} }+\frac{ 1 }{ 2^{2} }+\frac{ 1 }{ 2^{3} }+.....)\] \[\frac{ 1 }{ 3 }(\frac{ 1\times(1-\frac{ 1 }{ 2^{n} }) }{ 1-\frac{ 1 }{ 2 } })\] \[\frac{ 1 }{ 3 }\times \frac{ 1 }{ \frac{ 1 }{ 2 } }\] \[\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

Sum to infinite G.P

OpenStudy (anonymous):

\[S \infty = \frac{ a }{ 1-r }\]